OpenStudy (anonymous):

Rewrite the following expressions in equivalent y=mx + b form. a. 3x + 6y + 9 = 0 b. -2x + 6y - 4 = 0 c. -4x - 2y = 0

OpenStudy (anonymous):

a. y= -1/2x - 3/2 b. y= 1/3x + 2/3 c. y=-2x

OpenStudy (anonymous):

Thanks